OpenStudy (anonymous):

Please help: Solve for x: 5|2x − 2| + 8 = 18

OpenStudy (anonymous):

Part 1: Create an absolute value equation that has no solutions. Explain why this equation has no solutions. Part 2: Create an absolute value equation that has one solution. Explain why this equation has one solution. Part 3: Create an absolute value equation that has two solutions. Explain why this equation has two solutions.

OpenStudy (anonymous):

Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 (2 points): Use calories as the y−coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (anonymous):

thats all

OpenStudy (anonymous):

\[5|2x-2|+8=18\]ou have to isolate the absolute value (2x-2)\[5|2x-2|=10\]\[|2x-2|=2\]now you have to account for the outcome being both negative and positive like this\[2x-2=2~~~~~~~~~~2x-2=-2\]\[2x=4~~~~~~~~~~~~~~~~2x=0\]\[x=2~~~~~~~~or~~~~~~~x=0\]

OpenStudy (anonymous):

:) Thanks so much!!

OpenStudy (anonymous):

@cshalvey lend a hand? (:

OpenStudy (anonymous):

I need the bottom 2 questions only now.

OpenStudy (anonymous):

@cshalvey Can you help e? @yummydum reccommended you :)

OpenStudy (anonymous):

@jim_thompson5910 @experimentX @myininaya @physicsme

OpenStudy (experimentx):

Part 1: Create an absolute value equation that has no solutions. Explain why this equation has no solutions. |x| = -2 (no solutions) ... guess why??? Part 2: Create an absolute value equation that has one solution. Explain why this equation has one solution. |x-2| = 0 (one solution) ... guess why?? Part 3: Create an absolute value equation that has two solutions. Explain why this equation has two solutions. |x| = 2 (two solutions) ... guess why??

OpenStudy (experimentx):

Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 (2 points): Use calories as the y−coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? http://www.wolframalpha.com/input/?i=Plot+points+ {%2812%2C36%29%2C+%2818%2C+54%29} Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. use this formula: y - y1 = (y2 -y1)/(x2 - x1) * (x-x1) Part 3 (2 points): How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned. y - y1 = (y2 -y1)/(x2 - x1) * (35-x1) put ^^ 35 in this equation

OpenStudy (experimentx):

and your ((x1,y1),(x2, y2)) -> {(12,36), (18, 54)}